Medical command gives you the following orders for a rapid sequence intubation. You estimate the patient weighs at 175 pounds.

Succinylcholine is 200 mg/10ml.

Your order is 1.5 mg/kg.

What is the correct dose for this patient?
How many mL would you draw from this vial?
Answer #1

Ordered Succinylcholine= 1.5 mg/kg

Patient's weight= 175 pounds= 175/2.2kg= 79.54 kg

As per weight of the patient required dose= 79.545 X 1.5 mg=119.32 mg

Available Succinylcholine= 200mg/10 ml

= 20 mg/ml

For 119.32 mg required ml= 119.32/20

= 5.966 ml or almost 6 ml would be drawn from the vial